OpenStudy (anonymous):

given the area A of a circle, the formula r= √a/π can be used to calculate the radius of the circle. find the area when the radius is 9.9cm.

OpenStudy (jiteshmeghwal9):

put the values in the formula = \(\Large{r={\sqrt{a}\over\pi}}\) i.e.; \(\Large{{9.9={\sqrt{a}\over3.14}}}\)

OpenStudy (jiteshmeghwal9):

I use \(\Large{\pi=3.14}\)

OpenStudy (jiteshmeghwal9):

solve this :) can u do it ????

OpenStudy (anonymous):

do i move 9.9 on the other side?

OpenStudy (jiteshmeghwal9):

i mean solve for 'a'

OpenStudy (anonymous):

okay

OpenStudy (jiteshmeghwal9):

sorry no

OpenStudy (jiteshmeghwal9):

u must move 3.14 on left side

OpenStudy (jiteshmeghwal9):